# ode]python

### v2

@@ 3 Data generation
The data generation is carried out within the generate module. A single instance of the ThomasFermi class calculates quantities of interest for a single point in voltage-space only based on the device configuration specified in the physics module. To generate a complete CSD, a new simulation instance must be created for each pixel. However, since the gate voltages of neighboring pixels vary only slightly, it follows that the corresponding charge density $n(x)$ will also not change significantly between adjacent pixels. To optimize QDFlow performance, the result of the $n(x)$ calculation at one pixel is used as an initial condition when calculating $n(x)$ at adjacent pixels. This means that $n(x)$ must only be calculated from scratch once for each diagram.
QDFlow contains convenience functions for generating CSDs and rays in the generate module. Since the primary purpose of QDFlow is to generate data for training specialized ML models, it is essential that the resulting dataset captures the full range of variability observed in contemporary QD devices. To achieve this, QDFlow includes functionality to randomize nearly all physics parameters and to control the distributions from which each parameter is drawn. This capability is implemented via the PhysicsRandomization dataclass, which specifies each physics parameter as either a fixed value (when no randomization is desired, e.g., to allow regeneration of the same QD device), or a Distribution from which to draw the random values. A code example in Listing 3 shows how to import QDFlow and how to initialize a random configuration of physical parameters, with drawn from a distribution provided by the user.
@@
{code}
from qdflow import generate from qdflow.util import distribution
# Create a new dataclass instance that contains the default # randomization distributions for each physics parameter phys_rand = generate.PhysicsRandomization.default()
# Change the range from which mu can be drawn phys_rand.mu = distribution.Uniform(0, 1.2)
# Generate a list of 6 sets of random device parameters n_devices = 6 phys_params = generate.random_physics(phys_rand, n_devices)
Figure 3: Example code to generate a list of 6 randomized sets of device parameters. First, a PhysicsRandomization object is created, which defines the ranges and distributions, as appropriate, from which the physics parameters should be randomized. Distributions for each parameter can then be set as desired.
{code}
from qdflow import generate import numpy as np
# Create a set of physics parameters phys = generate.default_physics(n_dots=2)
# Set ranges (in mV) and resolution of plunger gate sweeps V_x = np.linspace(2., 16., 100) V_y = np.linspace(2., 16., 100)
# Generate a charge stability diagram csd = generate.calc_2d_csd(phys, V_x, V_y)
# Obtain the sensor readout in the form of a numpy array sensor_num = 0 sensor_readout = csd.sensor[:, :, sensor_num]
@@
Figure 5: Example code demonstrating how to generate a CSD from a PhysicsParameters object.
@@
This randomization framework, along with the dozens of configurable physics parameters, enables QDFlow to generate a highly diverse set of CSDs. Figure 4 shows six examples of CSDs generated with QDFlow. A snippet of code allowing to generate CSD data from a PhysicsParameters object is shown in Listing 5.
@@
While CSDs are extremely useful for visualizing charge states, they require extensive data collection. In practice, only a relatively small subset of points—the charge transitions—are of primary importance. As the number of dots grows, so does the dimensionality of the gate voltage space, rapidly making the exploration of the complete, multidimensional voltage space infeasible. This challenge is typically handled by measuring multiple 2D CSDs, each defined by a different pair of gates.
@@
In Ref. 11, an alternative method for assessing the charge state in QD devices, with 1D rays measured in multiple directions in the voltage space used in place of the 2D CSDs. This method greatly reduces the amount of data required for assessing the charge state of the device, but sacrifices some of the intuitive human interpretability provided by CSDs, necessitating the use of ML tools. To support the development of ML methods for ray-based analysis, QDFlow includes functionality for generating ray-based datasets, as shown in Listing 6.
{code}
from qdflow import generate import numpy as np from scipy.stats import qmc
# Create a set of physics parameters phys = generate.default_physics(n_dots=2)
# Generate quasirandom points inside a given area v_min, v_max = 2., 16. point_generator = qmc.Halton(d=2, scramble=False) initial_points = qmc.scale(point_generator.random(n=50), v_min, v_max)
@@
# Define a list of rays that will extend out from each point ray_length = 3. # length of rays in mV num_rays = 8 rays = ray_length * np.array([[np.cos(2*np.pi*i/num_rays), np.sin(2*np.pi*i/num_rays)] for i in range(num_rays)])
@@
# Generate ray data resolution = 100 # points per ray ray_data = generate.calc_rays(phys, initial_points, rays, resolution)
@@
Figure 6: Example code demonstrating how to generate ray data from a PhysicsParameters object.
@@ 4 Noise
To more faithfully emulate experimental conditions, QDFlow includes the noise module, which contains functionality for adding noise to both CSDs and rays, as well as several postprocessing functions designed to mimic effects of experimental measurements. The module implements several types of noise, including the white noise, pink ($1/f$) noise, telegraph noise, and latching effects, as well as stray transitions arising from nearby unintended dots [27, 23]. Postprocessing functions include adding gate-sensor coupling, adding a $\operatorname{sech}^{2}x$ blur, and adding Coulomb peak effects. Each noise in QDFlow can be controlled individually, with its magnitude defined relative to the scale of the CSD data, or as a predefined mixture. Similar to the physical parameter randomization discussed in Sec. 3, QDFlow supports designating the distributions from which each noise parameter is drawn via the NoiseRandomization dataclass.
@@
The simplest, white noise, is implemented by adding to each pixel a value drawn from a normal distribution with standard deviation given by the magnitude of the white noise. Pink noise is implemented by generating white noise in Fourier space with a uniform random phase and magnitude proportional to $1/\sqrt{f_{x}^{2}+f_{y}^{2}}$, where $f_{x}$ and $f_{y}$ are the components of each point in Fourier space, and then applying an inverse Fourier transform. This configuration allows for greater variability compared to telegraph noise. Alternatively, QDFlow also provides the option to add pink noise correlated along only the primary measurement axis, which corresponds to the more physical model of pink noise correlated in time.
Telegraph noise is applied along an axis corresponding to the primary measurement direction in experimental data. It consists of adding a value (drawn from a normal distribution with nonzero mean) to a line of several contiguous pixels. The length of the added line is randomly drawn from a geometric distribution. This allows the distribution of lengths of the telegraph noise to follow an exponential, as expected for two-level systems with finite excited-state lifetimes. This process is then repeated across the CSD, alternating the sign of the value added each time.
@@
The latching noise implemented in QDFlow can be controlled using one of two methods. The first and simpler legacy method to simulate latching is to shift each row of pixels by a random number of pixels drawn from a geometric distribution. This process produces a latching-like effect along the charge transitions; however, it is somewhat unrealistic since all transitions on the same row are shifted by the same amount, and because pixels far away from transitions are also displaced. A more physically realistic method relies on the nanowire simulation to calculate the sensor readout for both an excited charge state and a stable state at each pixel, similar to what is implemented in Ref. 17. The excited state chosen corresponds to the charge configuration most recently occupied prior to the most recent transition when sweeping the gate voltages. When latching noise is added, the sensor readout from the excited state replaces the stable-state readout for the first few pixels after each transition, with the number of pixels randomized each instance. In general, the second method is preferable as it more accurately reflects the experimental conditions; however, the first, legacy method is provided as a fallback when excited-state data are unavailable or computationally inconvenient, or impossible to obtain.
In addition to noise, several postprocessing functions can be applied to the data. A sensor-gate coupling in the form of a linear gradient along a random direction can be added. Convolving with a $\operatorname{sech}^{2}x$ kernel along the measurement axis introduces smoothing of the sharp transitions.
Finally, it is important to note that the physics simulation returns the value of the potential at each of the sensors; however, experimentally, the conductance of the sensor is the quantity that is measured. Therefore, we convert from potential to conductance by using a Coulomb peak lineshape of the form $G\propto\operatorname{sech}^{2}[A(V-V_{0})]$, where $G$ is the conductance of the sensor, $A$ is a parameter that determines the width of the Coulomb peak, $V$ is the potential at the sensor (the simulation output), and $V_{0}$ is the peak center [28].
@@
CSDs with a mixture of noises optimized for compatibility with experimental data are presented in Fig. 8(a)-(c). Figure 8(d) shows two rays with added noise. The exact amounts of each type of noise are randomized for each diagram. A code example showing how to generate noisy CSD data for a previously simulated sample CSD is shown in Listing 9.
{code}
from qdflow.physics import noise from qdflow.util import distribution import numpy as np
# Use data from previous example data = np.load("sensor_readout.npy")
# Create a new dataclass instance that contains the default # randomization distributions for each noise type noise_rand = noise.NoiseRandomization.default()
# How much noise to add, relative to the transition height upper bound noise_amount = 0.15
# Use a CorrelatedDistribution to randomize the white, pink, and # telegraph noise in such a way that the total always equals noise_amount num_dists = 3 dists = distribution.SphericallyCorrelated(num_dists, noise_amount).dependent_distributions() noise_rand.white_noise_magnitude = dists[0].abs() noise_rand.pink_noise_magnitude = dists[1].abs() noise_rand.telegraph_magnitude = dists[2].abs()
# Generate a random set of noise parameters noise_params = noise.random_noise_params(noise_rand)
# Add noise to the data noisy_data = noise.NoiseGenerator(noise_params).calc_noisy_map(data)
@@
Figure 9: Example code to add noise to a charge stability diagram. First, a NoiseRandomization object is created, which defines the ranges and distributions, as appropriate, from which the noise parameters should be randomized. A CorrelatedDistribution is used to randomize the different noise types while ensuring that the total noise amount is constant.
@@
Given the computational complexity of Thomas-Fermi calculations, the noise module is configured to assume that a complete noiseless CSD (or a ray-based data) has already been generated using the generate module. This approach gives us several advantages. First, it significantly reduces the computational overhead since multiple noise realizations with different relative noise strengths can be generated from a single noiseless CSD. Secondly, the modular approach adopted in QDFlow provides flexibility that cannot be achieved experimentally, where it is not possible to calibrate individual noise sources to the desired level. Additionally, it allows us to specify the magnitudes of each of the noise types relative to the local scale of the surrounding data points, which is important since the scale of the data points can vary across large CSDs. Finally, it allows for systematic benchmarking of ML algorithms since the noise level and type can be modified independently of the underlying physical configuration. For example, one can generate a single noiseless dataset and then apply different realizations of noise to study the robustness of a given algorithm under varying experimental conditions, as was done in Ref. 12.
@@ 5 Conclusion
Progress toward scalable quantum information technologies based on QD systems depends critically on overcoming the complexity of device operation and calibration with increasing number of QDs. Novel ML-based methods have emerged as powerful tools to address these challenges, but their effectiveness relies on access to large, diverse, and accurately labeled datasets. QDFlow was developed precisely to meet this need.
QDFlow differs from existing QD simulations in that it fully simulates the charge density function $n(x)$. By integrating a self-consistent Thomas-Fermi solver with a dynamic capacitance model, QDFlow provides a physics-informed simulation framework that goes beyond constant-capacitance approximations. This enables the generation of CSDs and ray-based data with features that evolve naturally with gate voltages, mimicking experimental behavior such as dot merging and transition slope variation.
@@
Looking forward, QDFlow represents a paradigm shift among quantum dot simulators. Whereas other simulations typically rely on constant-capacitance models that impose static couplings regardless of device state, QDFlow ties these parameters directly to the underlying physics through its self-consistent charge density, producing capacitances and observables that evolve dynamically with gate voltages. This distinction not only improves the connection with the experiment but also allows for the capture of nontrivial behaviors—such as QDs merging, fluctuating slopes, and disorder-induced effects—that are inaccessible to static-capacitance approaches.
As QD systems advance toward larger arrays and integration into functional quantum processors, the need for such realism will only grow. Possible future extensions of QDFlow, including multi-dimensional modeling, hybridization with experimental feedback loops, and systematic studies of robustness under different noise and disorder regimes, could establish it as a cornerstone for bridging theory, experiment, and ML in the quest for scalable quantum technologies. By releasing QDFlow as an open-source package, we aim to foster a shared foundation for accelerating progress in quantum dot technologies. We anticipate that the package will not only continue to advance automated QD control but also provide a flexible testbed for exploring broader questions at the intersection of condensed matter physics, quantum information, and machine learning.

### v3

@@ 3 Data generation
Data generation is performed within the generate module. A single instance of the ThomasFermi class calculates quantities of interest for a single point in voltage-space only based on the device configuration specified in the physics module. To generate a complete CSD, a new simulation instance must be created for each pixel. However, since the gate voltages of neighboring pixels vary only slightly, it follows that the corresponding charge density $n(x)$ will also not change significantly between adjacent pixels. To optimize QDFlow performance, the result of the $n(x)$ calculation at one pixel is used as an initial condition when calculating $n(x)$ at adjacent pixels. This means that $n(x)$ must only be calculated from scratch once for each diagram.
from qdflow import generate
from qdflow.util import distribution
# Create a new dataclass instance that contains the default
# randomization distributions for each physics parameter
phys_rand = generate.PhysicsRandomization.default()
# Change the range from which mu can be drawn
phys_rand.mu = distribution.Uniform(0, 1.2)
# Generate a list of 6 sets of random device parameters
n_devices = 6
phys_params = generate.random_physics(phys_rand, n_devices)
Listing 1: Example code to generate a list of 6 randomized sets of device parameters. First, a PhysicsRandomization object is created, which defines the ranges and distributions, as appropriate, from which the physics parameters should be randomized. Distributions for each parameter can then be set as desired.
QDFlow contains convenience functions for generating CSDs and rays in the generate module. Since the primary purpose of QDFlow is to generate data for training specialized ML models, it is essential that the resulting dataset captures the full range of variability observed in contemporary QD devices. To achieve this, QDFlow includes functionality to randomize nearly all physics parameters and to control the distributions from which each parameter is drawn. This capability is implemented via the PhysicsRandomization dataclass, which specifies each physics parameter as either a fixed value (when no randomization is desired, e.g., to allow regeneration of the same QD device), or a Distribution from which to draw the random values. A code example in Listing 1 shows how to import QDFlow and how to initialize a random configuration of physical parameters, with drawn from a distribution provided by the user.
@@
from qdflow import generate
import numpy as np
# Create a set of physics parameters
phys = generate.default_physics(n_dots=2)
# Set ranges (in mV) and resolution of plunger gate sweeps
V_x = np.linspace(2., 16., 100)
V_y = np.linspace(2., 16., 100)
# Generate a charge stability diagram
csd = generate.calc_2d_csd(phys, V_x, V_y)
# Obtain the sensor readout in the form of a numpy array
sensor_num = 0
sensor_readout = csd.sensor[:, :, sensor_num]
@@
Listing 2: Example code demonstrating how to generate a CSD from a PhysicsParameters object.
@@
This randomization framework, along with the dozens of configurable physics parameters, enables QDFlow to generate a highly diverse set of CSDs. Figure 3 shows six examples of CSDs generated with QDFlow. A snippet of code that generates CSD data from a PhysicsParameters object is shown in Listing 2.
@@
While CSDs are extremely useful for visualizing charge states, they require extensive data collection. In practice, only a relatively small subset of points—the charge transitions— is of primary importance. As the number of dots grows, so does the dimensionality of the gate voltage space, rapidly making the exploration of the complete, multidimensional voltage space infeasible. This challenge is typically handled by measuring multiple 2D CSDs, each defined by a different pair of gates.
@@
In Ref. [12], an alternative method for assessing the charge state in QD devices, with 1D rays measured in multiple directions in the voltage space used in place of the 2D CSDs. This method greatly reduces the data required to assess the device’s charge state but sacrifices some of the intuitive human interpretability provided by CSDs, necessitating the use of ML tools. To support the development of ML methods for ray-based analysis, QDFlow includes functionality for generating ray-based datasets, as shown in Listing 3.
from qdflow import generate
import numpy as np
from scipy.stats import qmc
# Create a set of physics parameters
phys = generate.default_physics(n_dots=2)
# Generate quasirandom points inside a given area
v_min, v_max = 2., 16.
point_generator = qmc.Halton(d=2, scramble=False)
initial_points = qmc.scale(point_generator.random(n=50), v_min, v_max)
@@
# Define a list of rays that will extend out from each point
ray_length = 3. # length of rays in mV
num_rays = 8
rays = ray_length * np.array([[np.cos(2*np.pi*i/num_rays),
np.sin(2*np.pi*i/num_rays)] for i in range(num_rays)])
# Generate ray data
resolution = 100 # points per ray
@@
ray_data = generate.calc_rays(phys, initial_points, rays, resolution)
@@
Listing 3: Example code demonstrating how to generate ray data from a PhysicsParameters object.
@@ 4 Noise
To more faithfully emulate experimental conditions, QDFlow includes the noise module, which contains functionality for adding noise to both CSDs and rays, as well as several postprocessing functions designed to mimic effects of experimental measurements. The module implements several types of noise, including white noise, pink ($1/f$) noise, telegraph noise, and latching effects, as well as stray transitions arising from nearby unintended dots [29, 25]. Postprocessing functions include adding gate-sensor coupling, adding a $\operatorname{sech}^{2}x$ blur, and adding Coulomb peak effects. Each noise in QDFlow can be controlled individually, with its magnitude defined relative to the scale of the CSD data, or as a predefined mixture. Similar to the physical parameter randomization discussed in Sec. 3, QDFlow supports designating the distributions from which each noise parameter is drawn via the NoiseRandomization dataclass.
@@
The simplest, white noise, is implemented by adding a value drawn from a normal distribution with standard deviation equal to the magnitude of the white noise to each pixel. Pink noise is implemented by generating white noise in Fourier space with a uniform random phase and magnitude proportional to $1/\sqrt{f_{x}^{2}+f_{y}^{2}}$, where $f_{x}$ and $f_{y}$ are the components of each point in Fourier space, and then applying an inverse Fourier transform. This configuration allows greater variability than telegraph noise. Alternatively, QDFlow also provides the option to add pink noise correlated along only the primary measurement axis, which corresponds to the more physical model of pink noise correlated in time.
Telegraph noise is applied along an axis corresponding to the primary measurement direction in experimental data. It consists of adding a value (drawn from a normal distribution with a nonzero mean) to a line of several contiguous pixels. The length of the added line is randomly drawn from a geometric distribution. This allows the distribution of telegraph noise lengths to follow an exponential distribution, as expected for two-level systems with finite excited-state lifetimes. This process is then repeated across the CSD, alternating the sign of the value added each time.
@@
The latching noise implemented in QDFlow can be controlled using one of two methods. The first, simpler legacy method to simulate latching is to shift each row of pixels by a random number of pixels drawn from a geometric distribution. This process produces a latching-like effect along the charge transitions; however, it is somewhat unrealistic because all transitions on the same row are shifted by the same amount, and pixels far from the transitions are also displaced. A more physically realistic method relies on the nanowire simulation to calculate the sensor readout for both an excited charge state and a stable state at each pixel, similar to what is implemented in Ref. [18]. The excited state chosen corresponds to the charge configuration most recently occupied prior to the most recent transition when sweeping the gate voltages. When latching noise is added, the sensor readout from the excited state replaces the stable-state readout for the first few pixels after each transition, with the number of pixels randomized each instance. In general, the second method is preferable, as it more accurately reflects the experimental conditions; however, the first, legacy method is provided as a fallback when excited-state data are unavailable, computationally inconvenient, or impossible to obtain.
In addition to introducing noise, several postprocessing functions can be applied to the data. A sensor-gate coupling in the form of a linear gradient along a random direction can be added. Convolving with a $\operatorname{sech}^{2}x$ kernel along the measurement axis introduces smoothing of the sharp transitions.
Finally, it is important to note that the physics simulation returns the value of the potential at each sensor; however, experimentally, the sensor’s conductance is measured. Therefore, we convert from potential to conductance by using a Coulomb peak lineshape of the form $G\propto\operatorname{sech}^{2}[A(V-V_{0})]$, where $G$ is the conductance of the sensor, $A$ is a parameter that determines the width of the Coulomb peak, $V$ is the potential at the sensor (the simulation output), and $V_{0}$ is the peak center [30].
@@
CSDs with a mixture of noises optimized for compatibility with experimental data are presented in Fig. 5(a)-(c). Figure 5(d) shows two rays with added noise. The exact amounts of each noise type are randomized for each diagram. A code example showing how to generate noisy CSD data for a previously simulated sample CSD is shown in Listing 4.
from qdflow.physics import noise
from qdflow.util import distribution
import numpy as np
# Use data from previous example
data = np.load("sensor_readout.npy")
# Create a new dataclass instance that contains the default
# randomization distributions for each noise type
noise_rand = noise.NoiseRandomization.default()
# How much noise to add, relative to the transition height upper bound
noise_amount = 0.15
# Use a CorrelatedDistribution to randomize the white, pink, and
# telegraph noise so that the total always equals noise_amount
num_dists = 3
dists = distribution.SphericallyCorrelated(num_dists,
noise_amount).dependent_distributions()
noise_rand.white_noise_magnitude = dists[0].abs()
noise_rand.pink_noise_magnitude = dists[1].abs()
noise_rand.telegraph_magnitude = dists[2].abs()
# Generate a random set of noise parameters
noise_params = noise.random_noise_params(noise_rand)
# Add noise to the data
noisy_data = noise.NoiseGenerator(noise_params).calc_noisy_map(data)
@@
Listing 4: Example code to add noise to a charge stability diagram. First, a NoiseRandomization object is created, which defines the ranges and distributions, as appropriate, from which the noise parameters should be randomized. A CorrelatedDistribution is used to randomize the different noise types while ensuring that the total noise amount is constant.
@@
Given the computational complexity of Thomas-Fermi calculations, the noise module is configured to assume that a complete noiseless CSD (or a ray-based data) has already been generated using the generate module. This approach gives us several advantages. First, it significantly reduces computational overhead, as multiple noise realizations with different relative noise strengths can be generated from a single noiseless CSD. Secondly, the modular approach adopted in QDFlow provides flexibility that cannot be achieved experimentally, where it is not possible to calibrate individual noise sources to the desired level. Additionally, it allows us to specify the magnitudes of each of the noise types relative to the local scale of the surrounding data points, which is important since the scale of the data points can vary across large CSDs. Finally, it allows for systematic benchmarking of ML algorithms, since the noise level and type can be independently adjusted without affecting the underlying physical configuration. For example, one can generate a single noiseless dataset and then apply different noise realizations to study the robustness of a given algorithm under varying experimental conditions, as was done in Ref. [13].
5 Benchmarking and limitations
To benchmark QDFlow, we find the mean runtime of the physics simulation for 10,000 randomly generated sets of physics parameters. Each simulation is run with at most 1,000 iterations and a relative convergence tolerance of $10^{-3}$. We are interested in comparing the contribution to the overall runtime from calculating $n(x)$ with the remaining portion of the simulation as the number of QDs increases. For each number of dots $N$, we obtain two runtimes: (i) the time required to compute the charge density profile $n(x)$, and (ii) the remaining time to construct the capacitance model, minimize Eq. (13), and evaluate the sensor output. The results for small- and mid-sized QD arrays are shown in Fig. 6(a) and Fig. 6(b), respectively.
For a small number of dots ($N\leq 20$), the main bottleneck is the initial part of the simulation where $n(x)$ is calculated. However, as the number of dots $N$ is increased, the minimization of Eq. (13) over integer charge configurations $\vec{Q}\in\mdmathbb{Z}^{N}$ for which the runtime scales exponentially in $N$ begins to dominate. In practice, we expect most use cases to involve no more than ten QDs, as long-range interactions tend to be negligible due to screening from the material and the nearby gates; thus, we focus attention on the cost of computing $n(x)$ in this regime.
Figure 6: Average runtime versus number of dots NN for (a) Thomas-Fermi solver and (b) capacitance model solver. Benchmarks were performed on a 2.8 GHz AMD Ryzen 5 7520U processor.
Because QDFlow calculates the capacitance model at each point in the CSD by explicitly solving for $n(x)$, it requires several orders of magnitude more runtime to generate CSDs than simulations based on constant capacitance models. Thus, for applications that require real-time data, faster simulations such as QDsim [17] or QArray [18] may be more suitable. On the other hand, for applications such as generating datasets for training and validation of ML models for the autonomous control of QD systems, the higher-fidelity data provided by QDFlow—including effects such as QDs merging across low barriers—often justifies the additional computational cost.
In addition to its relatively long runtime, QDFlow has several modeling limitations. While QDFlow is based on an underlying physical nanowire model, it is designed only to produce qualitatively realistic behavior. For example, gates are modeled as infinite cylinders arranged in a simple, idealized geometry, whereas in real devices, the shapes and arrangements of gates can be considerably more complex. Furthermore, the electron density $n(x)$ is calculated only in 1D, an approximation that can introduce errors for systems with non-negligible transverse extent. Despite being a 1D simulation, data generated with QDFlow is qualitatively similar to that from 2D QD arrays, provided that crosstalk is not too large. Influence from other dots can be included either during the simulation phase by extending the nanowire to include additional dots on the left and right sides, or during the noise-adding phase as unintended dot noise. However, the ability to model crosstalk from other dots remains somewhat limited, as QDFlow cannot reproduce data features from 2D architectures with tight couplings between nearby dots. Finally, the simulation uses a semiclassical approach and therefore does not include certain quantum effects, such as finite tunnel coupling effects and charge-state hybridization, which would be captured by more complex models such as the Hubbard model. Taken together, these approximations prevent QDFlow from generating quantitatively accurate data.
QDFlow also lacks certain capabilities present in other simulators. For example, it cannot currently model closed systems with a fixed number of charges isolated from the source and drain contacts—a functionality available in QArray [18]. Additionally, QDFlow does not currently model quantum effects due to finite tunnel coupling, which can cause broadening of the interdot transitions and rounding of sharp corners around the triple-points due to hybridization of charge states. For applications involving virtualization of gates, quantum effects are generally small enough that a semiclassical approximation is sufficient. However, for applications such as readout, where fine control of gates is needed, quantum effects are more important. If these features are relevant, QDarts contains a simulation of the finite tunnel coupling effect [19]. Finally, QDFlow is configured to model the state that minimizes the energy of the capacitance model, implicitly assuming that the system can transition freely between charge configurations. However, this assumption can break down for large arrays of QDs. For example, if a charge must transition between non-neighboring QDs passing through an intermediate QD that forms a large barrier, it may get “stuck” on one side even if it is energetically favorable to move to the other side. This will result in much more extreme latching effects than QDFlow can model.
QDFlow has been tested both modularly and holistically, through extensive unit testing and by comparing its output to experimental data. The QDFlow repository includes unit tests for every function in the package, ensuring full code coverage. The validity of QDFlow has been confirmed by the performance of multiple specialized ML models trained on QDFlow-generated data and deployed in real-world applications, as we discuss in Sec. 1 and in Sec. 6.
@@ 6 Conclusion
Progress toward scalable quantum information technologies based on QD systems depends critically on overcoming the complexity of device operation and calibration as the number of QDs increases. Novel ML-based methods have emerged as powerful tools to address these challenges, but their effectiveness relies on access to large, diverse, and accurately labeled datasets. QDFlow was developed precisely to meet this need.
QDFlow differs from existing QD simulations in that it fully simulates the charge density function $n(x)$. By integrating a self-consistent Thomas-Fermi solver with a dynamic capacitance model, QDFlow provides a physics-informed simulation framework that goes beyond constant-capacitance approximations. This enables the generation of CSDs and ray-based data with features that evolve naturally with gate voltages, mimicking experimental behavior, including dot merging and transition slope variations.
@@
QDFlow represents a paradigm shift among QD simulators. Whereas other simulations typically rely on constant-capacitance models that impose static couplings regardless of device state, QDFlow ties these parameters directly to the underlying physics through its self-consistent charge density, producing capacitances and observables that evolve dynamically with gate voltages. This distinction not only improves the connection to the experiment but also allows for capturing the nontrivial behaviors—such as QDs merging, fluctuating slopes, and disorder-induced effects—that are inaccessible to static-capacitance approaches.
As QD systems advance toward larger arrays and integration into functional quantum processors, the need for such realism will only grow. QDFlow package is currently being used to generate datasets for multiple ongoing research projects, and we intend to support it for the foreseeable future. Although the current release is stable, we anticipate adding further modules and functionality to the package. Possible future extensions of QDFlow, including multi-dimensional modeling, incorporation of additional quantum effects, hybridization with experimental feedback loops, and systematic studies of robustness across different noise and disorder regimes, could establish it as a cornerstone for bridging theory, experiment, and ML in the quest for scalable quantum technologies. We have designed the core class structure to allow seamless addition of new physical parameters and functions, and plan to update the generate module to support parallelization. By releasing QDFlow as an open-source package, we aim to foster a shared foundation for accelerating progress in quantum dot technologies. We anticipate that the package will not only continue to advance automated QD control but also provide a flexible testbed for exploring broader questions at the intersection of condensed matter physics, quantum information, and machine learning.